OpenStudy (anonymous):

2t + 3t + 5 = 2t - 3 Grade 9 work. please help me, i beg :(

OpenStudy (anonymous):

1) isolate the t's on one side. 2) then subtract and add as usual and find t as usual

OpenStudy (anonymous):

i would combine like terms on the left first : \[2t + 3t + 5 = 2t - 3 \] turns in to \[5t+5=2t-3\]

OpenStudy (anonymous):

then if you subtract \(2t\) from both sides you will get \[3t+5=-3\] subtract 5 to both sides , get \[3t=-8\] and finally divide both sides by 3 to get \(t=-\frac{8}{3}\)

OpenStudy (anonymous):

But Satellite how will we do this? Like your not explaining the (subtract 2t from both sides, subtract 5 to both and the divide both sides by 3)

OpenStudy (anonymous):

ok lets start at the end suppose i told you that \(3t=-8\) could you tell me what \(t\) was?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

ok lets try an easier one suppose i told you \(3t=12\) what would \(t\) be in this case?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

yes, and how do you know it? one reason is because \(3\times 4=12\) which is another way of saying \[4=\frac{12}{3}\]

OpenStudy (anonymous):

in other words, so solve \(3t=12\) you divide \(12\) by \(3\)

OpenStudy (anonymous):

2x + 5x = -9 + 4 4y 6y --- = ---- 2 3 32x + 2 = 4x +3 THESE QUESTIONS HELP

OpenStudy (anonymous):

now \(3\) does not go in to \(-8\) evenly, so we don't get a whole number as an answer, but you still divide \(-8\) by \(3\) but in this case you just write the answer as \(-\frac{8}{3}\) without actually dividing

OpenStudy (anonymous):

first one \[2x + 5x = -9 + 4 \] what is \(2x+5x\) ?

OpenStudy (anonymous):

Ok, now do the other ones. THANKS!

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

just do it please hurry :(

OpenStudy (anonymous):

do all the questions!

OpenStudy (anonymous):

if you just want the answers and are not worried about how to do them, don't ask here where there might be mistakes just do this: http://www.wolframalpha.com/input/?i=2x+%2B+5x+%3D+-9+%2B+4 it won't help you learn, but it will surely get all the answers right if that is what you are interested in

OpenStudy (anonymous):

i do want to learn

OpenStudy (anonymous):

that site isn't doing anything